OpenStudy (anonymous):

Mark had 1/2 cup sugar and borrowed 3/4 cup from Cary to add to his. How many cups did he have now

OpenStudy (anonymous):

First, we'll want to convert the 1/2 to fourths, so they have a common denominator, by multiplying top and bottom by 2. So, 1/2 is equal to 2/4. Now, simply add: \[2/4 + 3/4 = 5/4\] or 1 and 1/4 cup.

OpenStudy (anonymous):

1/2 + 3/4 = ? find common denominator 2/4 + 3/4 = 5/4 5/4 cups